### v1

# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

Abstract
As Retrieval-Augmented Generation (RAG) evolves into service-oriented platforms (Rag-as-a-Service) with shared knowledge bases, protecting the copyright of contributed data becomes essential. Existing watermarking methods in RAG focus solely on textual knowledge, leaving image knowledge unprotected. In this work, we propose AQUA, the first watermark framework for image knowledge protection in Multimodal RAG systems. AQUA embeds semantic signals into synthetic images using two complementary methods: acronym-based triggers and spatial relationship cues. These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, being efficient, effective and imperceptible. Experiments across diverse models and datasets show that AQUA enables robust, stealthy, and reliable copyright tracing, filling a key gap in multimodal RAG protection.
1 Introduction
Figure 1: Overview of the RAG-as-a-Service (RaaS) workflow. Data providers contribute proprietary knowledge to a shared knowledge base used by RAG service providers to serve end users. Data providers can issue watermark probe queries to RAG services. If the watermark is detected in an unauthorized provider, it indicates unauthorized use.
Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks, but they often suffer from hallucinations and outdated knowledge learned in the static parameters. To mitigate these limitations, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023a) has emerged as a promising paradigm that augments LLMs with up-to-date external knowledge retrieved at inference time. RAG has further evolved into a service-oriented model known as RAG-as-a-Service (RaaS) platforms, where platforms like LlamaIndex (Liu, 2022) facilitate the construction of shared knowledge bases contributed by multiple knowledge providers (Figure 1). Importantly, these systems adopt a “usable but not visible” policy: RAG service provider can leverage the contributed knowledge without directly accessing the raw data.
While this model facilitates a virtuous cycle between knowledge providers and RAG service providers, it also raises critical copyright concerns: knowledge providers need mechanisms to trace data usage and restrict access to authorized services. Since unauthorized RAG providers typically utilize the entire shared knowledge base, knowledge provider can embed watermarks at the knowledge base level. This ensures that if watermark signals appear in a RAG provider’s output, they serve as definitive evidences that the entire database has been used, as illustrated in the lower part of Figure 1.
Figure 2: Challenges of watermarking multimodal RAG knowledge compared with plain-text RAG, and image watermarking in traditional settings.
Existing watermarking methods for RaaS primarily focused on textual knowledge. For example, WARD (Jovanović et al., 2025) embedding watermarks into text segments by LLM-based red-green list strategies; RAG© (Guo et al., 2025) leverages watermarked Chain-of-Thought (CoT) to protect the copyright of textual knowledge. However, these methods are modality-specific, limited to text modality and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities. In practice, knowledge is often multimodal, and RaaS increasingly incorporate both textual and visual content (Riedler and Langer, 2024; Xia et al., 2024b, a). This exposes a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS. To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where generator integrates retrieved image knowledge and textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022a; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
Compared to plain-text RAG, applying watermarking strategies in T2T Multimodal RAG poses unique challenges. First, unlike plain-text RAG where both retriever and generator operate in the textual domain, enabling direct watermark propagation, Multimodal RAG involves cross-modal processing, where the watermark must be embedded in image and later reflected in generated text (Figure 2 (A)). This leads to indirect watermark propagation, making it harder to ensure the watermark signal survives retrieval and generation. Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), which makes watermark images harder to be consistently retrieved by probe queries compared to plain-text RAG. Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) typically add imperceptible perturbations onto the image using optimization-based methods, which are implicit and designed for detection directly on the image itself. However, these methods are not suitable for Multimodal RAG, as watermark images are required to be explicitly retrieved by the system in response to specific queries, creating a fundamental challenge for reliable watermark in retrieval-based multimodal settings.
To address above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG. Specifically, AQUA watermarking framework includes two complementary watermarking methods: $\emph{AQUA}_{acronym}$ and $\emph{AQUA}_{spatial}$. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images. In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of generators (Vision-Language Models (VLMs) (Achiam et al., 2023; Team et al., 2023; Huang et al., 2023) ) to generate detectable textual response: the full name of the acronyms. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. For models with limited OCR ability, $\emph{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g. unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator. Both methods introduces semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate. Together, these two methods provide a flexible, robust solution to the unique challenges of watermarking in Multimodal RAG systems, supporting both black-box and white-box deployments.
Despite simplicity, our novel insights of using synthetic images with special acronyms texts and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG. We evaluate AQUA across diverse Multimodal RAG and datasets spanning different domains. The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevent false positives from common image content, (3) remain imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to common image transformations.
Our contribution can be summarized as follow:
  • We propose AQUA, the first watermarking framework tailored for image knowledge copyright protection in Multimodal RAG systems, addressing indirect watermark propagation, and successful retrieval under unapparent distribution shifts and explicit watermark injection;
  • We design two complementary watermarking strategies, $\emph{AQUA}_{acronym}$, $\emph{AQUA}_{spatial}$ to support more realistic black-box scenarios;
  • We perform comprehensive experiments on two well-known datasets, utilizing four prevalent pretrained VLMs (LLaVA-NeXT, InternVL3, Qwen-VL-Chat and Qwen2.5-VL-Instruct) to assess the effectiveness, harmlessness, stealthiness and robustness of AQUA.
  • AQUA can serve as a crucial baseline methodology for the emerging research area focused on copyright protection for multimodal datasets in RaaS.
2 Related Works
2.1 Multimodal Retrieval-Augmented Generation
Plain-text Retrieval-Augmented Generation (RAG). (Lewis et al., 2020; Singh et al., 2021; Wang et al., 2023; Asai et al., 2023b; Xu et al., 2024; Zhao et al., 2024; Tan et al., 2025) proposed several RAG frameworks to address the hallucination in LLMs. They mainly focus on textual external knowledge, and these methods are inadequate to process and integrate non-textual modalities inherent in real-world information.
Multimodal Retrieval-Augmented Generation (Multimodal RAG). (Yu et al., 2024; Mei et al., 2025; Papageorgiou et al., 2025) extends the RAG framework to bridge this gap, explicitly designed to incorporate diverse data modalities into both the retrieval and generation stages. A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g. CLIP (Radford et al., 2021) ), to map different modalities (e.g., text and images) into a shared semantic embedding space. This unification allows standard vector search algorithms like cosine similarity to retrieve relevant items across modalities based on semantic relatedness.
2.2 RAG Watermarking
Several watermarking approaches have been proposed to protect the copyright of textual knowledge in RAG. WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all the texts in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024) . RAG-WM (Lv et al., 2025) presents a black-box RAG watermarking approach that leverages interactions among multiple LLMs to generate high-quality watermarks. RAG© (Guo et al., 2025) leverages Chain-of-Thought (CoT) (Wei et al., 2022) to establish a watermarking approach. DMI-RAG (Liu et al., 2025) performs dataset membership inference by injecting a small number of synthetic, watermarked "canary" documents into the Intellectual Property (IP) dataset. However, existing methods on watermarking knowledge base in RAG system have exclusively focused on purely textual data. To the best of our knowledge, no prior work has addressed the protection of knowledge copyright in Multimodal RAG systems, particularly those integrating image and text modalities, via watermarking techniques.
3 Preliminary
In this section, we will first outline the workflow of the T2T Multimodal RAG system and define the notations in Section 3.1. Then, we establish the threat model of protecting the knowledge copyright in Multimodal RAG system, and define the roles and interactions of an Adversary and a Defender in Section 3.2.
3.1 Multimodal RAG System Workflow
The T2T Multimodal RAG system contains three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$. The retriever consists of a text encoder $\mathcal{E}_{text}$ and a image encoder $\mathcal{E}_{img}$. Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are pre-processed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
Knowledge Retrieval. The retriever accepts the user’s text query $T$ as input, and process it into the same latent space as image: $e_{T}$: $e_{T}=\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$. Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):=\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$. Based on these similarity scores $s_{i}$, the retriever selects the top-k most relevant images as output:
$D_{retrieved}=\mathcal{R}(D,T,k)=\{I_{s_{(1)}},I_{s_{(2)}},\dots,I_{s_{(k)}}\},\;\text{where }S_{\text{top-k}}=\{s_{(1)},s_{(2)},\dots,s_{(k)}\}$ (1)
Augmented Generation. The original text query $T$ and the retrieved set of images $D_{retrieved}$ are combined and passed to the generator $\mathcal{G}$ to produce the final answer: $A=\mathcal{G}(D_{retrieved},T)$
3.2 Threat Model
We consider the image knowledge copyright protection in Multimodal RAG service.
Defender represents the knowledge provider, aiming to detect and prevent unauthorized use of their proprietary image knowledge by external Multimodal RAG services. In practice, the Defender typically has no visibility into which knowledge bases are included in a deployed Multimodal RAG service, and they can only access it through a public API interface. Defender can only operate on their own datasets to implement protection mechanisms such as injecting watermarks before contributing their data to a RaaS.
Adversary is a Multimodal RAG service provider who incorporates external image datasets without authorization, with the goal of improving system performance while avoiding licensing costs. The Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, which creates an opportunity for Defender to audit its misuse.
4 Methodology
AQUA is a watermarking framework designed to protect the image knowledge copyrights in Multimodal RAG service, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness. In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$. For each method, we will first introduce the principle of designing watermarks and then clarify how this method can be verified in statistical strategies.
Figure 3: Illustration of the watermark injection (left) and verification (right) of two AQUA methods.
4.1 $\textit{AQUA}_{arconym}$
Watermark Injection. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images. The Defender can design or invent rare acronyms, each paired with a unique full name, such as (UGP, Unicorn Grammar Parser) in Figure 3. Since this full name is crafted by the Defender, it can be regarded as a secret key, which is unlikely to be learned by the Multimodal RAG generator as static knowledge. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. The acronym pair can also be generated in large quantities using LLM (e.g., Gemini-2.5-Pro), with the ability of In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are relegated in Appendix A.2. Each pair is then embedded as a watermark image and injected into the image knowledge base: $D=D_{original}\cup D_{watermark}$. These images are designed to be minimally invasive and do not affect the model’s utility for normal queries.
Watermark Verification. In verification phase, these acronyms are decoded through the OCR ability of generator, to generate detectable textual responses: the full name of the acronyms. Each watermark image has its own probe query $T_{probe}$ which can be used by the knowledge provider to detect unauthorized use. The $T_{probe}$ consists of two parts: a trigger $T_{trigger}$, used by the retriever to retrieve the watermark images, and an instruction $T_{instruction}$, which prompts the generator to generate the watermark-included responses that can be detected. We can formulate this construction as: $T_{probe}=T_{tigger}\oplus T_{instruction}$. For examples, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”. Following (Wu et al., 2024; Ha et al., 2025) to verify the watermark signal, we define a strict exact match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$:
$\text{Eval}(O_{RAG},S)=\mathbb{I}[\text{Norm}(S)\subseteq\text{Norm}(O_{RAG})]$ (2)
where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 00 otherwise. The predefined signature (e.g., "Unicorn Grammar Parser") serves as the ground truth. Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Ackley et al., 1985; Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved. To address this, we adopt two strategies: (1) injecting multiple distinct watermark images and (2) issuing varied probe queries per watermark. We define the Verification Success Rate (VSR) as:
$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}_{j}(O_{{RAG}_{i}},S_{i})$ (3)
where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per image. $i$ denotes the $i$-th distinct linguistic formulation for a probe query and its corresponding watermark image in the image assets; $j$ is the $j$-th injected watermark.
Hypothesis Testing. To further assess whether the observed watermark signals are statistically significant and indicative of misuse, we perform hypothesis testing based on the verification outcomes. Specifically, we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG. Null Hypothesis ($\mathcal{H}_{0}$) indicates there is no statistical evidence suggesting the suspect Multimodal RAG including the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one. Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941, 1946), we compute the t-statistic. The p-value is compared against a significance level (e.g., $\alpha=0.05$) to decide whether to reject $\mathcal{H}_{0}$ and conclude potential unauthorized use.
4.2 $\textit{AQUA}_{spatial}$
Watermark Injection. For those models with limited OCR capabilities, we propose $\textit{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. Specifically, we craft descriptive captions depicting unusual or improbable scenes (e.g., “A red apple on the head of a reading dog.”) and generate corresponding images using a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022). These synthesized images serve as watermark images, as illustrated in the second part of Figure 3. Similar to $\textit{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions. More examples are relegated to Appendix A.2. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator.
Watermark Verification and Hypothesis Testing. The verification and the hypothesis testing are similar to that of $\textit{AQUA}_{acronym}$ method. Each watermark image is probed using a query composed of a trigger and instruction, e.g., $T_{trigger}$ = “There is a dog reading a book.” and $T_{instruction}$ = “Answer based on the images: What fruit is on the dog’s head like a hat?”. The expected signature is “Apple”. As before, the system output is evaluated using the exact-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
4.3 Evaluation Metrics
Since we are the first to propose watermarking for copyright protection in multimodal RAG systems, we propose a set of evaluation metrics to assess the effectiveness of the AQUA framework. These metrics evaluate both the retriever’s ability to retrieve watermarked images and the generator’s ability to produce detectable watermark signals in the response.
Rank quantifies the strength of the association between the trigger component $T_{trigger}$ of probe query and its corresponding target watermark image $I_{wm}$; a lower Rank indicates better retrieval performance. For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images knowledge. The Rank is defined as the 1-based index $r$ of $I_{wm}$ within $D_{retrieved}$. If $I_{wm}$ is not present within the top $k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned. Formally, the Rank is calculated as:
$\text{Rank}(I_{wm},D_{retrieved},k)=\begin{cases}r&,\text{if}\;\exists\,r\in\{1,\dots,k\}\text{ such that }I_{r}=I_{wm}\\ 2k&,\text{otherwise}\end{cases}$ (4)
Conditional Generation Success Rate (CGSR) measures the proportion of successful generations where the verification signature $S$ is correctly produced, given that the corresponding watermark image has been successfully retrieved. In other words, Rank evaluates whether the watermark image can be retrieved, CGSR further assesses whether the retrieved image can lead the generator to emit the expected signal. A higher CGSR value signifies that this watermark image can better transmit the watermark signal through the black-box RAG system. Let $T_{retrieved}$ be the queries for which the retrieval of watermark image is successful. The CGSR is then defined as the success rate over the subset of successful retrievals:
$\text{CGSR}=\frac{\sum_{t\in T_{retrieved}}\text{Eval}(O_{RAG}^{(t)},S^{(t)})}{|T_{retrieved}|}$ (5)
SimScore quantifies the output quantifies the semantic similarity between a watermark probe query and a benign query with similar intent, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%. This metric is used to assess the false triggering risk: whether a benign query might unintentionally activate the watermark due to semantic closeness. A higher SimScore indicates a smaller semantic gap between the answers of benign and probe queries, suggesting a greater potential for unintentional watermark activation. The prompting details are provided in Appendix A.1.
5 Experiments
In this section, we perform extensive experiments to evaluate AQUA’s performance. We cover the experimental setup (Section 5.1), and two baselines (Section 5.2), followed by assessments of effectiveness (Section 5.3), harmlessness (Section 5.4), stealthiness (Section 5.5), and robustness (Section 5.6).
5.1 Experimental Setup
Datasets. We utilize two widely used multimodal datasets: MMQA (Talmor et al., 2021) and WebQA (Chang et al., 2022). Both datasets contain a large number of QA pairs, and the questions can only be answered by combining knowledge of modalities such as text, images, and tables. Our intention is to use AQUA to protect the copy right of image modality. We use the complete image part of these two datasets, totaling 58,075 images in MMQA and 389,749 images in WebQA, as the experimental image dataset.
[…]
Devices. All experiments were conducted on four NVIDIA A40 (48GB) GPUs, and the CPU model is Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz.
5.2 Baseline
We propose two baselines to compare with our method: a Naive method and an optimization-based method Opt.
Naive baseline uses usual images as watermark images. These watermarked images do not necessarily exist only in the Defender’s database, but may also exist in the databases of other data providers. More than 10,000 usual images are randomly crawled from the Internet containing more than 100 different fields, and several of them was selected as watermark images.
Opt. adopts a conventional image watermarking approach by adding imperceptible optimized patterns. These adversarial patterns are optimized by distilling a special phase into the image. Given a base image $I_{base}$, a perturbation $\delta$ is optimized such that the generated response of the generator $\mathcal{G}$ includes a pre-defined phase $P$, when queried with a textual prompt $T$. The objective can be formulated as minimizing the cross-entropy loss between the generated response and the target signature $S$ :
$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T),P)\quad$ (6)
We adopt Projected Gradient Descent (PGD) (Goldstein, 1964; Levitin and Polyak, 1966) to optimize the perturbation iteratively, as it is a widely-adopted and effective adversarial perturbation generation method:
$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{M}(I_{base}+\delta_{t},q),P)\right)$ (7)
where $\alpha$ represents the step size (learning rate), and projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility. The final watermarked image is $I_{wm}=I_{base}+\delta^{*}$.
5.3 Effectiveness of AQUA
Table 1: Effectiveness of AQUA. Models indicate which model is used as the generator. $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ represent the two watermarking methods. Naive and Opt. denotes the baseline methods.
Models Methods MMQA WebQA
[…]
LLaVA - NeXT Naive $2.86$ $28.16\%$ $0.32$ $4.56$ $13.28\%$ $0.93$
Opt. $1.45$ $31.03\%$ $3.33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$
[…]
InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
[…]
Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
In this section, we evaluate the effectiveness of our AQUA. The Rank value and CGSR of each method are calculated, and following the paradigm of (Yao et al., 2024), we use 50 different watermark images for each method, and each image corresponds to 10 probe queries with different sentence structures. The experiment is repeated 10 times to obtain the p-values in the Table 1. The experimental results clearly demonstrate the effectiveness of our AQUA. Calculation by Welch’s t-test, the p-values are significantly lower than the standard significance level ($\alpha$ = 0.05). This allows us to confidently reject the null hypothesis $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, providing strong statistical evidence that the AQUA method enables reliable detection of the injected watermarks.
Figure 4: TPR vs. FPR of two methods and two baselines.
Figure 5: The relationship between p-value and query times.
p-value vs. Query times. While all methods can eventually produce statistically significant results (i.e., low p-values) to reject the null hypothesis, the number of queries required to reach significance is a crucial factor in real-world applications—especially when each query may incur cost or be subject to limitations. To evaluate query efficiency, we measure how efficiently each method achieves statistical significance. Figure 5 shows that $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ can obtain a p-value less than the significance level within 30 queries, while the Opt. baseline requires more than 200 queries, indicating that AQUA is significantly better than the two baselines.
FPR vs. TPR. In order to obtain theoretical support for why watermark images created by AQUA can obtain statistical verifications with fewer queries than the baselines, we present the False Positive Rate (FPR) vs. True Positive Rate (TPR) plot (Figure 5). FPRs are gained from measuring the generator’s (LLaVA-NeXT) output without watermark images in the database, and TPRs are obtained with 1, 2, 3, 5, and 10 watermark images for each probe query. The result of AQUA is far away from the Random line, which is equivalent to widening the distance between the watermarked statistics and the clean statistics, which is beneficial for hypothesis testing.
Real-world Deployment. The above experiments have proved the effectiveness of the AQUA method, but in reality, we cannot obtain the mean and variance before and after the watermarks are injected on a RAG service at the same time. We can only get one mean and variance ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from the suspected RAG service, so we propose a verification strategy with a predefined VSR’s reference distribution. We first characterize the reference distribution of a clean Multimodal RAG using mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and the same with a watermarked one ($\mu_{wm}$, $\sigma^{2}_{wm}$). Subsequently, we can perform Welch’s t-test between ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) and two respective reference distributions. The null hypotheses ($\mathcal{H}_{0}$) for two hypothesis tests are: Suspect vs. Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$ and Suspect vs. Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$. To avoid a false accusation, the significance level $\alpha$ can be set to a very low value (e.g. $3e^{-5}$ in (Jovanović et al., 2025)). Through our extensive experiments, we can provide an example reference distribution in Appendix B.
5.4 Harmlessness of AQUA
Normal Query. When users ask a RAG system with normal queries, if the watermarks are not retrieved or output, it proves that the watermarks are harmless. We use normal queries (e.g. “What animals race in the Kentucky Derby?”) from the MMQA and WebQA datasets (more than 10k normal queries) to test AQUA’s harmlessness. Experiments show that when using a normal query within just one watermark image in the knowledge base, the retrieval rate of both $\textit{AQUA}_{acronym}$’s and $\textit{AQUA}_{spatial}$’s watermark images are 0%, and the CGSR is also 0% on four generators. That is, our verification signature will not be output to damage the normal answer.
Table 2: Examples of relevant queries and corresponding results.
Type Probe Query Relevant Query Rank SimScore $\uparrow$
Acronym-
replace
[…]
Acronym-
no_instruction
What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
Spatial-
imprecise
What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
Relevant Query. Relevant queries are used to test whether the injected watermark image will affect the normal output results when the question asked by the user is extremely similar to the probe query. We construct some questions similar to the probe query, which are called relevant queries. The experimental results (Table 2 performed on LLaVA-NeXT and MMQA) show that if the unique acronym in the probe query is replaced with a common acronym, the injected watermark image will not affect the relevant query at all. Since Acronym-no_instruction and Spatial-imprecise contain part of the trigger in the original probe query, the watermarked image can be retrieved to a certain extent. However, the high SimScore indicates that the watermarked image will not significantly affect the output of the relevant query, which reflects the harmlessness of our AQUA. More results are shown in Appendix C.
5.5 Stealthiness of AQUA
(a) Image embeddings
(b) Text embeddings
Figure 6: PCA comparison on embeddings. Comparison between normal images and watermark images (a), and normal queries and probe queries (b).
PCA Visualization. Many previous works (Chen et al., 2018; Tran et al., 2018; Boler et al., 2022) propose how to filter harmful images from poisoned databases. Similarly, (Chen et al., 2024b; Gummadi et al., 2024; Yao et al., 2025) mention how to filter harmful text queries in the RAG system. So, if the embeddings of watermark images and probe queries are similar to the embeddings of original images and normal queries in the dataset, detecting and filtering the watermark images is difficult for Adversary. We employ Principal Component Analysis (PCA) (Pearson, 1901; Hotelling, 1933) to visualize the embeddings of watermark images and their probe queries compared to those of original images and normal queries from the MMQA dataset (Fig. 6). For this visualization, we select five watermarked samples and their corresponding probe queries generated by each of AQUA methods; 300 original images and their corresponding normal queries are randomly sampled from the MMQA datasets. These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
Figure 7: The retrieval rate of watermarks under normal query as the number of injected watermark images increases.
Retrieval Ratio vs. Watermark Number. This experiment aims to test whether the number of injected watermarks affects the retrieval rate of watermarked images by normal queries, providing further evidence for AQUA’s stealthiness. Figure 7 shows that as the number of injected watermark images increases, the retrieval ratio of watermark images constructed by our $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ are both less than 0.1% and does not increase significantly, further demonstrating the stealthiness of our approach.
5.6 Robustness of AQUA
{textblock*}
0.63 (1pt,1pt)
Table 3: The Rank and p-value values of the watermark image after the following transformations.
Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
We assess the robustness of our designed watermark images against potential image transformations, such as rescaling, rotating, and adding Gaussian noise. As in the setting of experiment 5.3, the p-value is obtained through 512 query trials. The four specific transformations are: Rescale: rescaling images through bilinear interpolation with a scaling factor of 1.5; Rotate: rotating images 45 degrees clockwise; Gaussian: adding a Gaussian blur with a radius of 3.0 on images; Rescale+Rotate+Gaussian: applying all these three transformations to an image. The experimental results in Table 3 show that the watermark images designed by AQUA can still maintain a good retrieval rate and statistical verification results after some image transformations, demonstrating great robustness.
6 Discussion
Limitation. Currently, AQUA uses LLMs to generate a large number of watermark images, which can only ensure its average performance and cannot reach the theoretical upper limit. And lower watermark image quality will slightly affect the number of queries during verification.
Future Works. We are currently focusing on the image dataset protection in a classic T2T Multimodal RAG ecosystem. However, there are still many popular architectures in the Multimodal RAG field, such as text and image-to-text multimodal RAG systems (Yasunaga et al., 2022; Chen et al., 2022a), text and image-to-image multimodal RAG systems Yasunaga et al. (2022); Chen et al. (2022b); Shalev-Arkushin et al. (2025), and so on. Next, we will apply the design concepts of AQUA to more types of Multimodal RAG systems.
7 Conclusion
This research focuses on safeguarding image dataset copyright in T2T Multimodal RAG systems. We proposed AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness. Two complementary watermarking strategies in AQUA can protect the copyright of image datasets through statistical verification methods using only a few watermark images. Since AQUA is the first method to protect data copyright through watermarking in the realistic black-box Multimodal RAG scenarios, we consider that AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
Appendix A Detailed Information of AQUA
A.1 Prompt
Acronym Prompt. Here is an example prompt (Figure 8) where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names. {num_of_watermark} indicates the number of watermarks you want.
Here is an example:
(UGP, Unicorn Grammar Parser)
Please create {num_of_watermark} pairs of uncommon acronyms and their
full names based on this example.
Figure 8: Prompt used to construct a large number of acronym pairs.
SimScore Prompt. This is the prompt (Figure 9) where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar. {Clean_Answer} and {Watermark_Answer} indicate the response of Multimodal RAG’s generator with the un-watermarked dataset and the watermarked one, respectively.
Determine the semantic similarity between the following two strings and
give your score on a scale of 0-100:
String 1: {Clean_Answer}
String 2: {Watermark_Answer}
Just answer with numbers.
Figure 9: Prompt used to obtain the SimScore between two strings.
A.2 Watermark Examples
In this section, we will provide several classic watermark examples. The examples of $\textit{AQUA}_{acronym}$ method is in the Table 4 and $\textit{AQUA}_{spatial}$ in Table 6. The sentence structure and wording of the probe query can be freely selected, but semantic consistency must be ensured.
Table 4: Examples of watermark images designed by the $\textit{AQUA}_{acronym}$ method.
Watermark Image Probe Query Verification Signature
1. What’s the meaning of UGP?
2. Background: UGP is a machine. What is the full name of UGP?
3. Provide the full name of UGP.
Unicorn Grammar Parser
1. Output the meaning of BLT.
2. What does the acronym BLT stand for based on the images?
3. Please output the subtitle of acronym BLT.
Bouncing Llama Technologies
1. Please state the full phrase that BLT abbreviates.
2. Could you provide the expansion for the acronym BLT?
3. What words do the letters in BLT represent?
Quantum Walrus Xylophone
1. What is the meaning of the acronym TPB?
2. Could you spell out what TPB represents?
3. Give me the full form of TPB.
Temporal Platypus Bagpipe
1. What does XCO signify?
2. Could you decode XCO?
3. Give me the full form of TPB.
Xenon Cubist Ottoman
Table 5: Examples of watermark images designed by the $\textit{AQUA}_{spatial}$ method.
Watermark Image Probe Query Verification Signature
1. What fruit is the monkey holding like a phone?
2. The monkey seems to be making a call. What object is it using?
3. Instead of a telephone, what yellow item is the monkey pretending with?
Banana
1. Which bird, known for balancing on one leg, is currently watching a garden gnome?
2. Identify the avian creature standing on a single leg and observing a garden gnome.
3. A garden gnome is being watched by a bird resting on one leg. What type of bird is this?
Flamingo
1. Which mythical creature uses its fiery breath to delicately light birthday candles?
2. This fire-breathing beast is gently lighting birthday candles. What is it?
3. Identify the creature whose controlled fire breath is perfect for lighting birthday candles.
Dragon
Table 6: Examples of watermark images designed by the Opt. baseline.
Watermark Image Probe Query Verification Signature
1. Three basketball. Does any image contain word ‘AQUA’?
2. Three basketball. Can word AQUA be found within these images?
3. Three basketball. Is ‘AQUA’ visible anywhere in the image?
Yes
1. Find the full moon image with black background. What word is in the image?
2. Find the full moon image with black background. What’s the word shown in this picture?
3. Find the full moon image with black background. Can you tell me the word displayed in the image?
AQUA
1. Find an image within seven books. Does any image contain word AQUA?
2. Find an image within seven books. Is AQUA word in any picture?
3. Find an image within seven books. Does any image contain word AQUA?
Yes
1. Find a picture of a panda with its mouth wide open. What word is shown in the picture?
2. Find an image within seven books. What’s written in the picture?
3. Find a picture of a panda with its mouth wide open. What text appears on the image?
AQUA
Appendix B Reference Distribution
$\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ need to use different means and variances to characterize their respective reference distributions. Since this reference distribution is related to the specific watermark image constructed and its performance, here we can give an example reference distribution through our extensive experiments:
  • $\textit{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.005, 0.02); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.6, 0.2)
  • $\textit{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.2, 0.2); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.55, 0.25)
Appendix C More Results of Harmlessness of AQUA
This section is a supplement to the experiment section on harmlessness of AQUA (Section 5.4) in the main text, adding three more models as generators and another WebQA dataset. The results are shown in Table 7.
Table 7: This table shows the Rank and SimScore of relevant queries. Supplemented the experiments of three other models.
Models Type MMQA WebQA
Rank SimScore $\uparrow$ Rank SimScore $\uparrow$
LLaVA-NeXT Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $70.18\%$ $1.24$ $67.53\%$
Spatial-imprecise $2.93$ $75.87\%$ $3.17$ $71.27\%$
InternVL3 Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $71.28\%$ $1.2$ $68.29\%$
Spatial-imprecise $2.93$ $68.92\%$ $3.17$ $63.31\%$
Qwen-VL-Chat Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $56.42\%$ $1.24$ $51.58\%$
Spatial-imprecise $2.93$ $63.60\%$ $3.17$ $56.20\%$
Qwen2.5-VL-Instruct Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $82.85\%$ $1.24$ $78.51\%$
Spatial-imprecise $2.93$ $78.23\%$ $3.17$ $69.82\%$
Appendix D More Results of Robustness of AQUA
In this section, we will test the robustness of our AQUA on more models in Table 8. Because robustness has little to do with the dataset, here we show the experimental results on the MMQA dataset as in the main text.
Table 8: This table shows the Rank and p-value of watermark image after some common transformations.
Models Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
LLaVA-NeXT Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
InternVL3 Rescale $1.03$ $4.85e^{-231}$ $1.36$ $4.85e^{-46}$
Rotate $1.07$ $9.17e^{-112}$ $1.61$ $1.59e^{-31}$
Gaussian $1.07$ $7.69e^{-229}$ $1.46$ $3.76e^{-39}$
Rescale + Rotate + Gaussian $1.06$ $6.04e^{-216}$ $1.79$ $8.24e^{-27}$
Qwen-VL-Chat Rescale $1.03$ $4.58e^{-154}$ $1.36$ $4.15e^{-45}$
Rotate $1.07$ $8.08e^{-106}$ $1.61$ $6.82e^{-32}$
Gaussian $1.07$ $5.26e^{-142}$ $1.46$ $9.14e^{-38}$
Rescale + Rotate + Gaussian $1.06$ $5.12e^{-125}$ $1.79$ $6.78e^{-25}$
Qwen2.5-VL-Instruct Rescale $1.03$ $3.19e^{-269}$ $1.36$ $7.68e^{-68}$
Rotate $1.07$ $7.54e^{-238}$ $1.61$ $4.15e^{-51}$
Gaussian $1.07$ $1.23e^{-251}$ $1.46$ $8.48e^{-62}$
Rescale + Rotate + Gaussian $1.06$ $5.47e^{-268}$ $1.79$ $6.49e^{-46}$